# Can We Catch the Elephant? The Evolvement of Hallucination Evaluation on Natural Language Generation: A Survey

### v1

@@ 3.3 Datasets and Benchmarks
In addition to the evaluation models above, other studies have proposed helpful datasets and benchmarks to measure hallucination across tasks and to assess the efficacy of evaluators.
While building models, some works also release annotated datasets as well. For the summarization task,  Maynez et al. (2020) find severe extrinsic hallucination in the reference of XSum dataset, and label XSumFaith with hallucination type (intrinsic or extrinsic) and hallucination span position. CoGenSumm (Falke et al., 2019) is collected from the validation set of CNN/DM and is used to rerank generated summaries. From both CNN/DM and XSum, QAGS (Wang et al., 2020) annotates each sentence with a binary choice based on whether the sentence is faithful to the article. However, the binary factuality of texts can be difficult to determine. FRANK (Pagnoni et al., 2021) collect annotation based on a more fine-grained defined typology of factual errors. Polytope (Huang et al., 2020) provides both SF and WF annotations to measure both extractive and abstractive summarization.  Devaraj et al. (2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang & Lapata, 2017) for text simplification. In dialogue generation, Dialogue NLI (Welleck et al., 2018) provides three-type labels for NLI models of the entailment of sentence pairs.
Given the large amount of evaluation metrics proposed, some benchmarks are designed to reevaluate these metrics. SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao & Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection. Wizard of Wikipedia (Dinan et al., 2018), CI-ToD (Qin et al., 2021), BEGIN (Dziri et al., 2022b), FaithDial (Dziri et al., 2022a) and TopicalChat (Gopalakrishnan et al., 2023) can help measure consistency evaluators on dialogue generation. Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom & Hill, 2021) can evaluate metrics across a series of text generation tasks.
@@ 4 Evaluation Methods After LLMs
With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in the natural language as human experts when evaluating. For SF and WF in LLM scenarios, it is mainly presented in the form of context consistency or factual imagination. Some works use LLMs to evaluate specific tasks, while other works evaluate the hallucination level of LLMs themselves. In this section, we will elaborate on both aspects.
4.1 LLMs as Tool
Recent LLMs are powerful tools that can be prompted by instructions for evaluation. The criteria can be designed by task, so it is more flexible and can evaluate from multiple aspects.
@@ 4.1 LLMs as Tool
Following the works of Sec. 3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses LLM as an NLI model to generate scores regarding to the prompts and generated texts. It breaks long source documents into chunks of arbitrary length to measure the supportive level of the source text span and also provides a long-form dialogue dataset ScreenEval from TV scripts and summaries pulled from the SummScreen.  Chen et al. (2023b) leverage prompting methods to evaluate summarization task. It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios and validates on a large range of summarization datasets.
@@
Expanding to a larger scope of tasks, GPTScore (Fu et al., 2023) and G-Eval (Liu et al., 2023) both provide multi-faceted evaluation frameworks including consistency. GPTScore (Fu et al., 2023) covers a range of tasks including dialogue response generation, summarization, data-to-text, and NMT. When evaluating different tasks, normally, LLMs will be fed with the task introduction and the evaluation criteria as a prompt. GPTScore uses LLMs till GPT-3 (Brown et al., 2020) and can customize evaluation for different tasks. G-Eval (Liu et al., 2023) is backboned with GPT-4, using CoT and a form-filling paradigm. When evaluating, it is asked to generate a CoT of evaluation steps formatted as a form. Then the final score is the summation of weighted scores with predefined weight. The experiment results on summarization and dialogue generation showed that CoT can help GPT-4 better understand the evaluation criteria. Other than these two works, Wang et al. (2023b) studies the feasibility of ChatGPT as a good evaluator for three tasks. This paper shows us a way to build evaluators based on ChatGPT with or without reference. It explores both direct assessment and one-to-five stars ranking processes. ChainPoll (Friel & Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
Using LLMs for consistency evaluation generally results in better human correlation, which is an important direction for the development of evaluators. Human evaluation can be expensive, especially the experts who hold domain specialities. However, the performance of the above tools still is sensitive to the LLM’s capacity and the design of the prompts.
4.2 LLMs as Object
LLMs have a unique form, which themselves are dialogue systems and also can involve NLG tasks in the conversation. To implement evaluation, due to the flexible form and nonuniform test datasets, many works focus on proposing their own datasets, evaluation indicators, and establishing benchmarks. Among them, some studies assess LLM’s grasp of general facts and also updated world knowledge. Additionally, measuring hallucinations in long-context/generation and reasoning scenarios is also challenging, as is assessing LLM’s performance in specific domains and non-English contexts. Here we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from these aspects.
General Fact:
In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool and UFO both show competitive evaluation frameworks for hallucination. FacTool (Chern et al., 2023) is a tool-augmented framework that includes five main components: claim extraction, query generation, tool querying, evidence collection, and agreement verification. It includes multiple tasks and multiple domains, which makes it a relatively holistic evaluation method. UFO (Huang et al., 2024) provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge. The process has three modules: fact unit extraction, fact source verification, and fact consistency discrimination. It extracts entity-centric question-answer pairs as fact units and checks if the fact is consistent with the source. CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess. It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue, as intrinsic evaluation.
Utilizing LLM’s inherent capabilities for consistency checking is also a viable approach. SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge. It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency. SelfCheckGPT sentence-level score is computed by incorporating previous metrics, such as BertScore, question answer, and LLM prompt. Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling. The cross-checking period includes question-level and model-level. By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers.
Benefiting from the open source of recent LLMs, it is possible to access the parameters and dense distribution. EigenScore (Chen et al., 2024) exploits dense semantic information of LLMs to measure semantic consistency and diversity. Based on the observation of the existing extreme features in the LLM penultimate layer, it also designs a method (Test Time Feature Clipping) to prevent LLMs generate overconfident hallucinations.
@@ 4.2 LLMs as Object
In addition to the aforementioned evaluation frameworks, many benchmarks have also been proposed to facilitate a more comprehensive evaluation of LLMs. TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough. It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse. This paper suggests that, for improving truthfulness, only scaling up models is less promising than finetuning using training objectives, and larger models were more informative despite being less truthful. HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text. Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations. FACTOR (Muhlgay et al., 2023) is a benchmark that can measure LLM factuality concerning a controlled set of facts in a generation setting, including three sub-benchmarks: Wiki-FACTOR, News-FACTOR and Expert-FACTOR. It follows the error types from FRANK (Pagnoni et al., 2021) performs a multi-choice factual evaluation task.
@@
Furthermore, for evaluating the metrics themselves, RealHall (Friel & Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF. FELM (Zhao et al., 2024) also provides a meta-evaluation for metrics. It broadens the conventional understanding of world knowledge factuality to diverse domain data including world knowledge, science, math, recommendation, and reasoning. It uses ChatGPT to generate the responses for prompts and do segmentation, while humans are asked to label factuality and error reasons, error types, and references for each sample.
Fresh Fact:
@@
The world is constantly changing. How can we determine whether LLMs possess dynamic knowledge? The following benchmarks concentrate on constructing time-sensitive datasets to enable the evaluation of LLMs’ capacity to incorporate up-to-date information. FreshQA (Vu et al., 2023) is a QA benchmark that can question current world knowledge and also with false premises to the LLMs. The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions, which test the LLMs from different aspects. KoLA (Yu et al., 2023) use both Wikipedia and continuously collected emerging news and novels to evaluate the world knowledge of LLMs. It develops a self-contrast metric by contrasting two completions generated by the same model and gets the similarity score. RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system. It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated. ERBench (Oh et al., 2024) verified facts through entity-relationship (ER) model. It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
Long Context/Generation:
@@
Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge. FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs. It breaks the generated text into a series of fine-grained atomic facts by InstructGPT and then each atomic fact is assigned a binary label. The data is about biographies of people sampled from English Wikipedia. Rather than considering whether the fact is globally truthful, FACTSCORE refers to a given source of knowledge. In this way, it can avoid information conflictions between different sources. BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination. It uses an entailment model on two synthetic datasets named SenHallu and AbsHallu to check the hallucination level of LLMs. RV (Yang et al., 2023) performs zero-resource hallucination detection by reverse validation. It focuses on passage-level detection rather than sentence-level, which holds greater practical significance. This method first constructs a question for the entity of user concern from LLM’s output, then uses it to query the LLM and compare the answers. This paper also publishes a benchmark PHD for the evaluation of passage-level detection methods.
Factual Reasoning:
To evaluate the factual reasoning abilities of LLMs, Laban et al. (2023) build a benchmark SUMMEDITS. This work provides a three-step protocol for inconsistency detection benchmark creation and implements it in a 10-domain benchmark. It shows current LLMs are not yet proficient at complex factual reasoning
Domain Specific:
@@
Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications. Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT). Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
Chinese Language:
@@
Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem. ChineseFactEval (Wang et al., 2023a) conducts the hallucination evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, including general knowledge, scientific, medical, legal, financial, mathematics, and modern Chinese history scenarios. UHGEval (Liang et al., 2023) provides a Chinese-language dataset for LLM hallucination evaluation, generated by Chinese LLMs in news domain. The evaluation method uses LLMs for discriminative evaluation (binary classification), selective evaluation (multiple-choice question), and generative evaluation (use existing lexical and semantic metrics). HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors. Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture to create a new dataset. It also instructs GPT-4 to act as a judger to get the consistency rate. The result shows this benchmark maintains a challenging benchmark while 18 out of 24 LLMs rate below 50%.
@@
It should be noted that there is a small portion of overlap between Sec. 4.2 and Sec. 4.1, meaning that some approaches utilize LLMs to evaluate LLMs (Chern et al., 2023; Zhao et al., 2024) as well. For instance, due to its outstanding performance, GPT-4 can do annotation instead of human sometimes (Cheng et al., 2023). In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect. This is because LLMs enable open-domain QA and can even leverage search engines for retrieval. Thus, assessing fresh facts is also an important aspect. Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse. Factual errors can also accumulate in reasoning or multi-hop question answering. Furthermore, in other application domains such as the medical field and non-English language, some preliminary explorations have been conducted, which call for more work participation.

### v2

@@ 3.3 Datasets and Benchmarks
Alongside evaluation models, datasets and benchmarks have been proposed to measure hallucination across tasks and assess evaluator efficacy.
For summarization,  Maynez et al. (2020) find severe extrinsic hallucination in the reference of XSum dataset, and label XSumFaith with hallucination type (intrinsic or extrinsic) and hallucination span position. CoGenSumm (Falke et al., 2019) is collected from the validation set of CNN/DM and is used to rerank generated summaries. From both CNN/DM and XSum, QAGS (Wang et al., 2020) annotates each sentence with a binary choice based on whether the sentence is faithful to the article. However, the binary factuality of texts can be difficult to determine. FRANK (Pagnoni et al., 2021) collect annotation based on a more fine-grained defined typology of factual errors. Polytope (Huang et al., 2020) provides both SF and WF annotations to measure both extractive and abstractive summarization. Falsesum (Utama et al., 2022) mimics intrinsic errors and extrinsic errors to augment summary datasets for NLI models.  Devaraj et al. (2022) label 3 types of factual errors for data collected from Newsela (Xu et al., 2015) and Wikilarge (Zhang and Lapata, 2017) for text simplification. In dialogue generation, Dialogue NLI (Welleck et al., 2018) provides three-type labels for NLI models of the entailment of sentence pairs.
Given the large amount of evaluation metrics proposed, some benchmarks are designed to reevaluate these metrics. SummEval (Fabbri et al., 2021), SummaC (Laban et al., 2022), Dialsummeval (Gao and Wan, 2022), and AGGREFACT (Tang et al., 2023) focus on summarization hallucination evaluation or detection. Wizard of Wikipedia (Dinan et al., 2018), CI-ToD (Qin et al., 2021), BEGIN (Dziri et al., 2022b), FaithDial (Dziri et al., 2022a) and TopicalChat (Gopalakrishnan et al., 2023) can help measure consistency evaluators on dialogue generation. Beyond a single task, TRUE (Honovich et al., 2022) and BEAMetrics (Scialom and Hill, 2021) can evaluate metrics across a series of text generation tasks.
@@ 4 Evaluation Methods After LLMs
With the rapid development of LLMs and their human-like performance, LLMs possess instruction-tuned LLMs, like ChatGPT (OpenAI, 2022) or GPT-4 (Achiam et al., 2023), which offer an advantage that can take instruction in natural language as human experts when evaluating. For SF and WF in LLM scenarios, they are mainly presented in the form of context consistency or factual imagination. Some works prompt LLMs as direct evaluators, while other works design a pipeline with multiple steps to do the assessment. In this section, we will elaborate on both aspects.
4.1 Direct Evaluation
Recent LLMs are powerful tools that can be prompted by instructions for text evaluation. The criteria can be designed according to different tasks, so it is more flexible and can assess from multiple aspects. Meanwhile, to evaluate the level of LLMs’ hallucination, a variety of methods have been developed as well. This section explores these methods using direct assessment. These approaches include prompting the LLMs to generate responses or using a classifier to decide whether the input contains hallucinations.
4.1.1 LLMs as tool
In this section, we introduce the methods that leverage LLMs as tools to conduct hallucination evaluation.
@@ 4.1.1 LLMs as tool
Following the works of Sec. 3, SCALE (Lattimer et al., 2023) focuses on long-form dialogue and uses the LLM as an NLI model to generate scores. It breaks long source documents into chunks to measure the supportive level of the source text span.  Chen et al. (2023b) leverage prompting methods to evaluate summarization task. It designs chain-of-thoughts (CoT) and sentence-by-sentence prompts on few-shot and zero-shot scenarios. TrueTeacher (Gekhman et al., 2023) elaborates a data augmentation method, which uses LLM to label the generated summaries from CNN/DM, as well as multilingual data.
@@
Expanding to a larger scope of tasks, GPTScore (Fu et al., 2023) and G-Eval (Liu et al., 2023) both provide multi-faceted evaluation frameworks including consistency. GPTScore (Fu et al., 2023) covers a range of tasks including dialogue, summarization, data-to-text, and NMT. When evaluating different tasks, normally, LLMs will be fed with the task introduction and the evaluation criteria as a prompt. G-Eval (Liu et al., 2023) uses CoT and a form-filling paradigm. When evaluating, it is asked to generate a CoT of evaluation steps formatted as a form. It shows that CoT can help GPT-4 better understand the evaluation criteria. Other than these two works, Wang et al. (2023b) study the feasibility of ChatGPT as a good evaluator for three tasks. This paper shows us a way to build evaluators based on ChatGPT with or without reference. ChainPoll (Friel and Sanyal, 2023) also uses ChatGPT to build a hallucination detection method with the CoT paradigm and considers it to be a binary classification problem.
Using LLMs for evaluation is simple and convenient. Despite the increasing trust in LLMs as they grow in size and capability, determining how to apply LLMs effectively to ensure their stability and reliability remains an open question.
4.1.2 LLMs as Object
In this section, we present recent evaluators and benchmarks for evaluating the hallucination level of LLMs from many significant aspects.
@@ 4.1.2 LLMs as Object
General Fact: TruthfulQA (Lin et al., 2022) is a benchmark to measure whether a language model is truthful enough. It brings up three concerns of the hallucination in the QA system: accidental misuse, blocking positive applications, and malicious misuse, and hedging is an acceptable solution for LLMs to be truthful in this work. HaluEval (Li et al., 2023) finds that ChatGPT is likely to generate unverifiable hallucinations and hard to identify the hallucinations in the generated text. Experiments suggest that extra knowledge and reasoning steps for LLMs can help alleviate hallucinations. FACTOR (Muhlgay et al., 2023) can measure LLM factuality concerning a controlled set of facts in a generation setting. It follows the error types from FRANK (Pagnoni et al., 2021) and performs a multi-choice factual evaluation task. Benefiting from the open source of recent LLMs, it is possible to access the parameters and dense distribution. EigenScore (Chen et al., 2024) exploits dense semantic information of LLMs to measure semantic consistency and diversity. Based on the observation of the existing extreme features in the LLM penultimate layer, it also use Test Time Feature Clipping to prevent LLMs generate overconfident hallucinations.
@@
Furthermore, for evaluating the metrics themselves, RealHall (Friel and Sanyal, 2023) is a benchmark for evaluation methods and contains both closed- and open-domain hallucinations, which correspond to SF and WF. FELM (Zhao et al., 2024) also provides a meta-evaluation for metrics. It broadens the conventional understanding of world knowledge factuality to diverse domain data including world knowledge, science, math, recommendation, and reasoning. It uses ChatGPT to generate the responses for prompts and do segmentation, while humans are asked to label factuality and error reasons, error types, and references for each sample.
@@
Fresh Fact: The world is constantly changing. How can we determine whether LLMs possess dynamic knowledge? The following benchmarks concentrate on constructing time-sensitive datasets to enable the evaluation of LLMs’ capacity to incorporate up-to-date information. FreshQA (Vu et al., 2023) is a QA benchmark that includes questions about current events and also inputs with false premises to the LLMs. The questions are divided into four types according to the freshness referring to the ever-changing world, including never-, slow-, fast-changing, and false-premise questions. RealTimeQA (Kasai et al., 2024) designs an annotation framework and submission workflow to support the updated evaluation system. It tests on both open- and closed-book QA systems and shows the importance of QA systems to be keep-dated. ERBench (Oh et al., 2024) verifies facts through entity-relationship (ER) model. It can construct questions using templates from updated ER databases and evaluates LLMs with binary or multi-choice QA tasks.
@@
Long Context/Generation: Despite recent advancements in LLMs enabling them to better handle long texts, hallucination evaluation in long context or generation remains a challenge. FACTSCORE (Min et al., 2023) focuses on long-form text generated by LLMs. It breaks the generated text into a series of fine-grained atomic facts and then each atomic fact is assigned a binary label. The data is about biographies of people sampled from English Wikipedia. It refers to a given source of knowledge, which avoids information conflictions between different sources, but challenges the comprehensiveness and accuracy of the knowledge data. BAMBOO (Dong et al., 2023) provides a multi-task long context benchmark for LLMs which also can detect hallucination, which uses an entailment model on two synthetic datasets.
@@
Domain Specific: Hallucinations generated by LLMs in the medical field can have serious consequences in healthcare applications. Therefore MedHalt (Pal et al., 2023) designs a pipeline to implement three fact-related tests: False Confidence Test (FCT), None of the Above Test (Nota), and Fake Questions Test (FQT). Through these tests, MedHalt can effectively identify and filter hallucinations within the medical domain.
@@
Chinese Language: Numerous Chinese LLMs have also emerged along with the trend and hallucination is also a crucial problem. ChineseFactEval (Wang et al., 2023a) conducts the evaluation in seven scenarios, covering areas where generative AI may be applied in daily life, and specifically includes the modern Chinese history scenario. UHGEval (Liang et al., 2023) provides a Chinese-language hallucination dataset, generated by Chinese LLMs in news domain. The evaluation method uses LLMs for discriminative, selective and generative evaluation. HalluQA (Cheng et al., 2023) is a benchmark categorizing hallucinations into imitative falsehoods and factual errors. Inspired by TruthfulQA, it summarizes the question patterns and combines them with Chinese culture.
@@
It should be noted that there is a small portion of overlap between Sec. 4.1.2 and Sec. 4.1.1, meaning that some approaches utilize LLMs to evaluate LLMs (Zhao et al., 2024) as well. For instance, due to its outstanding performance, GPT-4 can sometimes do annotation instead of human (Cheng et al., 2023). In contrast to evaluators before LLMs that primarily focus on SF, methods outlined in this section not only can address SF evaluation but also pay more attention to the WF (general fact) aspect, according to Figure 1. This is because LLMs enable open-domain QA and can even leverage search engines for retrieval. Thus, assessing fresh facts is also an important aspect. Though the latest LLMs can support long-text processing, challenges still persist in analyzing long contexts or generating long outputs, leading to hallucinations caused by long-distance discourse. How to catch the discourse-level errors while maintaining the ever-changing context is still challenging. Furthermore, in other application domains such as the medical field and non-English language, the collection of domain-specific data and the design of methodologies require further investigation, particularly in unique contexts such as medical statistics, legal terminology, and historical culture.
4.2 Pipeline Evaluation
Apart from utilizing LLMs for direct assessment, some evaluation methods do not rely on specific knowledge datasets or the pre-existing knowledge of LLMs. Instead, they leverage relevant evidence acquired through LLMs or search engines, benefitting from their enhanced search capabilities. Additionally, some methods involve the model generating answers multiple times to cross-check consistency to get the assessment outcome.
In a pipeline of fact extraction, evidence confirmation and fact consistency verification, FacTool (Chern et al., 2023) and UFO (Huang et al., 2024) both show competitive evaluation frameworks for hallucination. FacTool is a tool-augmented framework that includes includes multiple tasks and domains, which makes it relatively holistic. UFO provides a factuality verification method with multiple fact sources including human-written evidence, reference documents, search engine results, and LLM knowledge. CONNER (Chen et al., 2023a) also relies on external evidence to test LLMs the world knowledge they possess. It performs knowledge-intensitive tasks, including open-domain question answering and knowledge-grounded dialogue.
Utilizing LLM’s inherent capabilities for consistency checking is another viable approach. SelfCheckGPT (Manakul et al., 2023b) is motivated by the simple idea that LLM tends to have self-consistency if it possesses certain knowledge. It uses the scheme of black-box zero-resource hallucination detection and compares multiple sampled responses then measuring consistency. Rather than self-consistency, ${}\mathbf{SAC^{3}}$ (Zhang et al., 2023a) performs cross-checking hallucination detection based on sampling. By leveraging semantically equivalent question perturbation and cross-model response consistency checking, it can effectively identify the consistency and factuality of the answers. KoLA (Yu et al., 2023) use both Wikipedia and continuously collected emerging news and novels to evaluate the world knowledge of LLMs. It develops a self-contrast metric by contrasting two completions generated by the same model and gets the similarity score. RV (Yang et al., 2023) performs zero-resource hallucination detection by reverse validation. It focuses on passage-level detection rather than sentence-level, which holds greater practical significance. To evaluate the factual reasoning abilities of LLMs, Laban et al. (2023) build a benchmark SUMMEDITS. This work provides a three-step protocol for inconsistency detection benchmark creation and implements it in a 10-domain benchmark. It shows current LLMs are not yet proficient at complex factual reasoning.
These methods employ LLMs as tools to assess the hallucination level of models, including the LLMs themselves. While using multi-step approaches for evaluation offers strong interpretability, it is prone to error accumulation throughout extraction, reasoning, or multi-hop question answering. Similarly, having the model generate multiple answers or using multiple models to generate answers relies heavily on the models’ confidence and robustness. This dependency can lead to the oversight of high-confidence errors. We additionally present four case examples in Appendix D, covering both SF errors and WF errors. Table 2 shows that perfect evaluation for different types of hallucination errors remains challenging, and the evaluators also possess varying degrees of transferability.